OpenStudy (anonymous):

express the area under the curve y=1/x from 2 to 4 as a limit with sample points chosen to be the midpoints of the corresponding segments. Do not evaluate the limit.

OpenStudy (fibonaccichick666):

First, Draw a picture.

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry its crappy

OpenStudy (anonymous):

\[required area=\int\limits_{2}^{4}y dx=\int\limits_{2}^{4}\frac{ 1 }{ x }dx=2 \to 4 \left[ \ln x \right]=\ln 4-\ln 2=\ln \frac{ 4 }{ 2 }=\ln 2\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks, but wat about the midpoints?

OpenStudy (anonymous):

it is the area under the curve and between x=2 and x=4

OpenStudy (anonymous):

got it thnks.

OpenStudy (anonymous):

you are welcome.